OpenStudy (anonymous):

Imagine you are teaching a fellow student how to solve: 3 − 5(x + 2) = 3(x + 1) + 6 In your own words, explain the process for solving this equation. Please include your work at each step of the process along with the final answer.

OpenStudy (anonymous):

need help

OpenStudy (jiteshmeghwal9):

3-5x+10=3x+3+6 =>-5x+13=3x+9 =>-5x-3x=9-13 =>-8x=-4 =>x=-4/-8=1/2.

OpenStudy (anonymous):

thx

OpenStudy (jiteshmeghwal9):

yw:D